D Motion Begle Date: 1/20/2019 1:00:00 PM-De Date: 1/30 (10%)Problem8: Ascagull flies at a velocity of 9.5 m/s relative tothe air straight into thc wind. 1 2019 9.0000 AM End Daies 3hao S% Part (a) If is takes the bird 23 min to travel S.8 km relative to the Earth, what is the velocity of the wind, is meters per secoed, elative to the Earth? Take the directios of travel of the seaguill to be positive Grade Summary cotan cos Partial °Degrees Radians per hist Hin ▲ 50% ib) If the bird Dams around and flies with the wind. how long, n minnes, will it take toreturn 5.8 km? MacBook Pro 4 0 0
Answer #1

A)

Vs = 5800/(23 x 60) = 4.20 m/s

Velocity of air relative to ground, VAG = 4.2 - 9.5 = - 5.30 m/s

B)

Vs = - 9.5 - 5.3 = - 14.80 m/s

Time, t = 5800/14.80 = 391.97 sec = 6 min 31 sec

Comment in case any doubt please rate my answer .....